Find the coordinates of the y-intercept and x-intercept, in that order.
4y-7=3x

Answer by rfer(16322)

4y-7=3x
4y=3x+7
y=3/4x+7/4
-----
y int =7/4
(0,1 3/4)
------------
y=3/4x+7/4
3/4x=-7/4
x=-7/4*4/3
x=-28/12
x=-2 1/3
x int =-2 1/3
(-2 1/3,0)